OpenStudy (anonymous):

Find the mean for the given set of data. -4, -3, -1, -1, 0, 1 is it -3/4?

OpenStudy (anonymous):

close! -4/3

OpenStudy (anonymous):

-1 1/3?

OpenStudy (amistre64):

-10+1 = -9/6

OpenStudy (anonymous):

yeah, -1 and -1/3

OpenStudy (amistre64):

read a -2 in there ..... -9+1 = -8, -8/6 ... so toniks good

OpenStudy (anonymous):

it should be -8/6, not -9/6

OpenStudy (anonymous):

<3